OpenStudy (anonymous):

If cosθ = 0.3, sinθ is approximately: Select one: a. 0.98 b. 0.95 c. 0.91 d. 0.70

OpenStudy (anonymous):

@chosenmatt do you know what it is :0

OpenStudy (chosenmatt):

no sorry but if i had to guess i would choose A

OpenStudy (anonymous):

@eliassaab

OpenStudy (aaronq):

find the angle first, then plug into the sine function. \(\theta=Acos(0.3)\)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

You should fan me. ;)

OpenStudy (anonymous):

\[ \sin (\theta )=\sqrt{1-\cos ^2(\theta )}=\sqrt{1-0.3^2}=0.954 \] It is B